# Hölder regularity of harmonic functions on metric measure spaces

### v2

@@ 4 Poisson equation on unbounded MMD spaces
$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m+\sum_{k\leq k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{k\leq k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
@@
The idea is to consider the heat equation “$\Delta(P_{t}f)=\frac{\partial}{\partial t}(P_{t}f)$” for fixed $t$ as a Poisson equation.
@@
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,32\Psi^{-1}(t))}|P_{t}f|\mathrm{d}m$
$+\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}$
$+\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}.$
@@
$\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,32\Psi^{-1}(t))}|P_{t}f|\mathrm{d}m$
$\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,32\Psi^{-1}(t))}\lVert P_{t}f\rVert_{L^{\infty}(X;m)}\mathrm{d}m$
@@
For the second term, by UHK($\Psi$) and [23, THEOREM 4], we have the following estimate of the time derivative of the heat kernel
@@
is bounded by some constant depending only on $C_{2}$ and $\Phi$, see also [3, Page 944, Line -5], [20, Equation (2.5)], hence
@@
$\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}$
$\lesssim\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(x,2^{k})}\left(\frac{\lVert f\rVert_{L^{\infty}(X;m)}}{t}\right)^{p}\mathrm{d}m\right)^{1/p}$
@@
$\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert f\rVert_{L^{\infty}(X;m)}.$
@@
By UHK($\Psi$) and [23, THEOREM 4], we have the following estimates of the heat kernel
@@
Take any $t\in(0,+\infty)$, $x,y_{1},y_{2}\in X$. We now distinguish the following four cases.
@@
For the case (i), we obtain the result directly from UHK($\Psi$) as follows.
@@
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}p_{t}(x,z)m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left|\frac{\partial}{\partial t}p_{t}(x,z)\right|^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left|\frac{\partial}{\partial t}p_{t}(x,z)\right|^{p}m(\mathrm{d}z)\right)^{1/p}.$ (6.1)
For the case (ii), we have $d(x,y_{1})<128\Psi^{-1}(t)$ and $d(x,y_{2})<128\Psi^{-1}(t)$. By Equation (6.1), we have
@@
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
@@
For the case (iii), we have $d(x,y_{1})\geq 128\Psi^{-1}(t)$. Then in Equation (6.1), for any $z\in B(y_{1},32\Psi^{-1}(t))$, we have
@@
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$\leq\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
@@
For the case (iv), similar to the case (iii), we also have
@@
Lemma 6.1.
([25, Lemma 5.2]) Let $A\in(0,+\infty)$. Then there exists $C\in(0,+\infty)$ depending only on $A$, $\beta_{1}$, $\beta_{2}$ such that

### v3

@@ 4 Poisson equation on unbounded MMD spaces
$\lesssim\frac{\Phi(r)}{\Psi(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m+\sum_{k\leq k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}+\sum_{k\leq k_{1}}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,2^{k})}|f|^{p}\mathrm{d}m\right)^{1/p}.$
@@
The idea is to consider the heat equation $\Delta(P_{t}f)=\frac{\partial}{\partial t}(P_{t}f)$” for fixed $t$ as a Poisson equation.
@@
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,32\Psi^{-1}(t))}|P_{t}f|\mathrm{d}m$
$+\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}$
$+\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}.$
@@
$\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,32\Psi^{-1}(t))}|P_{t}f|\mathrm{d}m$
$\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,32\Psi^{-1}(t))}\lVert{P_{t}f}\rVert_{L^{\infty}(X;m)}\mathrm{d}m$
@@
For the second term, by UHK($\Psi$) and [24, THEOREM 4], we have the following estimate of the time derivative of the heat kernel
@@
is bounded by some constant depending only on $C_{2}$ and $\Phi$, see also [4, Page 944, Line -5], [21, Equation (2.5)], hence
@@
$\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}$
$\lesssim\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,2^{k})}\left(\frac{\lVert{f}\rVert_{L^{\infty}(X;m)}}{t}\right)^{p}\mathrm{d}m\right)^{1/p}$
@@
$\sum_{k\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{k})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y,2^{k})}|\frac{\partial}{\partial t}(P_{t}f)|^{p}\mathrm{d}m\right)^{1/p}\lesssim\frac{\Phi(\Psi^{-1}(t))}{\Psi(\Psi^{-1}(t))}\lVert{f}\rVert_{L^{\infty}(X;m)}.$
@@
By UHK($\Psi$) and [24, THEOREM 4], we have the following estimates of the heat kernel
@@
Take any $t\in(0,+\infty)$, $x,y_{1},y_{2}\in X$. We now distinguish the following four cases:
@@
For case (i), we obtain the result directly from UHK($\Psi$) as follows.
@@
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}p_{t}(x,z)m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},2^{j})}\left|\frac{\partial}{\partial t}p_{t}(x,z)\right|^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{2},2^{j})}\left|\frac{\partial}{\partial t}p_{t}(x,z)\right|^{p}m(\mathrm{d}z)\right)^{1/p}.$ (6.1)
For case (ii), we have $d(x,y_{1})<128\Psi^{-1}(t)$ and $d(x,y_{2})<128\Psi^{-1}(t)$. By Equation (6.1), we have
@@
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
@@
For case (iii), we have $d(x,y_{1})\geq 128\Psi^{-1}(t)$. Then in Equation (6.1), for any $z\in B(y_{1},32\Psi^{-1}(t))$, we have
@@
$\lesssim\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(C_{2}d(x,z),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$\leq\frac{\Phi(16\Psi^{-1}(t))}{\Psi(16\Psi^{-1}(t))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},32\Psi^{-1}(t))}\frac{C_{1}}{V(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)m(\mathrm{d}z)$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{1},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
$+\sum_{j\leq[\log_{2}16\Psi^{-1}(t)]}\Phi(2^{j})\left(\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(y_{2},2^{j})}\left(\frac{C_{1}}{tV(x,\Psi^{-1}(t))}\exp\left(-\Upsilon(\frac{C_{2}}{2}d(x,y_{1}),t)\right)\right)^{p}m(\mathrm{d}z)\right)^{1/p}$
@@
For case (iv), similar to case (iii), we also have
@@
Lemma 6.1 ([26, Lemma 5.2]).
Let $A\in(0,+\infty)$. Then there exists $C\in(0,+\infty)$ depending only on $A$, $\beta_{1}$, $\beta_{2}$ such that